OpenStudy (anonymous):

Would some one please help? Given that (7,-6) is on the graph of f(x), find the corresponding point for the function f(x+2).

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

solve \[x+2=7\] for \(x\)

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

takes one step, what do you get?

OpenStudy (anonymous):

ok so if \(f(x)\) has the point \((7,-6)\) then \(f(x+2)\) contains the point \((5,-6)\) because \[f(5+2)=f(7)=-6\]

OpenStudy (anonymous):

so the answer would be (7,-6)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-6 ?

OpenStudy (anonymous):

i think i wrote the answer directly above

OpenStudy (anonymous):

oh (5,-6)

OpenStudy (anonymous):

would you help me with another? Given that (2,9) is on the graph of f(x) find the corresponding point for the function f(x+2)

OpenStudy (anonymous):

@satellite73